# thanks: Corresponding authorthanks: Corresponding author

### v3

@@ I Introduction
The deep learning landscape is still rapidly evolving, with newer architectures such as Transformers Vaswani (2017); Khan et al. (2022); Lin et al. (2022) and denoising diffusion probabilistic models (DDPMs) Ho, Jain, and Abbeel (2020); Yang and Sommer (2023, 2023)-along with scalable diffusion models integrated with Transformers (DiT) Peebles and Xie (2023), showing superior performance in tasks like image generation. This evolution raises an important question: Can the latest advancements in machine learning enable real-time, accurate predictions of flow fields, effectively addressing the challenges of computational efficiency and accuracy?
@@
Efficient and accurate prediction of high Reynolds number airflow around airfoils is crucial for advancements in aerodynamics research and the aviation industry, as it directly impacts airfoil optimization, optimal flow control, and the design of more efficient aerodynamic systems. The application of deep learning in aerodynamics has evolved progressively from foundational flow field prediction to efficient optimization design. Early research (Sekar, 2019) established a speed foundation for replacing traditional CFD through a hybrid CNN-MLP architecture (where CNNs extract geometric features and MLPs map flow parameters), enabling real-time flow field prediction, though limited to steady laminar flows. To address data dependency issues, Wu (2021) introduced a Generative Adversarial Network (daGAN), leveraging a two-phase training mechanism (cGAN pre-training followed by dual-generator fine-tuning) to overcome sparse data constraints and enhance generalization under small-sample conditions. Li (2020) developed Deep Learning-based Geometric Filtering (DLGF), utilizing autoencoder-based compression of the design space to reduce variable dimensionality from hundreds to tens, thereby circumventing the curse of dimensionality and accelerating optimization workflows. Within this framework, Wang (2022) integrated a CNN-based Reduced Order Model (ROM) with a fusion strategy of pooling dimensionality reduction and radial basis functions to handle nonlinear flow optimization under large disturbances. Concurrently, Li (2021) addressed low-Reynolds-number separation bubbles by integrating the XFOIL solver with a transition-sensitive turbulence model, achieving efficient airfoil design for UAVs through tailored modal decomposition.
@@ II.2 Diffusion Transformer
In recent years, Transformers have achieved remarkable success across various domains, including natural language processing and computer vision Vaswani (2017). Meanwhile, diffusion models, as a class of generative models, have shown significant potential in image generation tasks. The Diffusion Transformer (DiT) combines the strengths of Transformers with a diffusion-based architecture, demonstrating outstanding performance in diverse applications Peebles and Xie (2023). Despite these advances, the application of DiT models to aerodynamic simulation remains unexplored. To the best of our knowledge, this study represents the first application of a DiT-based model trained on Reynolds-averaged Navier-Stokes (RANS) simulation data, with a particular focus on predicting flow around airfoils. In addition, in line with prior DiT-based work Ovadia et al. (2023), our study addresses the challenge of real-time inference. A detailed description of the methodology is provided in the following section.
@@ III.4 Forward Diffusion Process
In the forward diffusion process, we followed the canonical procedures described in literature Ho, Jain, and Abbeel (2020); Sohl-Dickstein et al. (2015). As shown in Figure 2, we gradually add noise to the data $\bm{x}_{0}$ (the true flow field) over several time steps given the input conditions $\bm{\Psi}$, transforming it into a noisy sample $\bm{x}_{T}$. The noise schedule $\beta_{t}$ is pre-defined and fixed, increasing with each timestep. At each timestep $t$, the conditional probability for adding noise is given by:
@@ III.5 Reverse Denoising Process and the Loss
The reverse denoising process $q(\bm{x}_{t-1}|\bm{x}_{t},\bm{\Psi})$ follows a Gaussian distribution, with its mean and variance derived asHo, Jain, and Abbeel (2020):
@@ III.6 Physics Informed Loss
where $\sigma_{c}$, $\sigma_{p}$, and $\sigma_{v}$ denote the mean absolute values of the convective, pressure, and viscous terms, respectively, used for normalization. Here $\nabla^{2}u_{x}$ and $\nabla^{2}u_{y}$ are computed component-wise (vector Laplacian), and $\mathbf{g}=(g_{x},g_{y})$ represents the body force per unit mass.
@@ IV.1.1 Datasets and Data preparation
In this study, we utilize three datasets and a test set, as described in Thuerey et al. (2020). These datasets are generated using RANS simulations to compute the velocity and pressure distributions of incompressible flows around airfoils. The simulations cover a wide range of Reynolds numbers ($Re=[0.5,5]$ million) and angles of attack ($\pm 22.5^{\circ}$), providing a diverse set of aerodynamic conditions. A total of 1505 unique airfoil shapes were sourced from the UIUC database and combined with randomly sampled freestream conditions to create the training inputs. The RANS solutions were computed using the Spalart-Allmaras (SA) turbulence model, a widely used one-equation model for industrial applications. The computations were carried out with OpenFOAM, an open-source CFD software.
@@
The test set comprises 90 images, which do not appear in the training data and remain unseen during all stages prior to the testing phase. The Reynolds number distribution in the test set spans a wide range, from $2.04\times 10^{6}$ to $2.00\times 10^{7}$], encompassing 30 different airfoil types. All the datasets used in this study are publicly available TUM-I Repository 1 (2018); TUM-I Repository 2 (2018).
@@ IV.1.2 Configuration
We employ the DiT (Scalable Diffusion Models with Transformers Peebles and Xie (2023)) model as the backbone for AeroDiT and adopt the same encoder and decoder models as in Ho, Jain, and Abbeel (2020). We use three configurations: DiT-B/2, DiT-L/2, and DiT-XL/2. The suffix "/2" indicates that the model operates on a latent feature grid that is downsampled by a factor of 2 in spatial resolution, achieved through a patch size of $2\times 2$ during tokenization.
@@
After training, the conditional encoder is retained as the encoder, and the spatial decoder is retained as the decoder. The trained encoder and decoder are then used to train the DiT model for 2,500,000 iterations. The model is trained using a global batch size of 4 for 150 epochs with the AdamW optimizer (learning rate of $10^{-4}$, no weight decay). Unless otherwise specified, the sampling process during inference uses 100 steps.
@@ IV.2 Results and Discussion
We provide more specific details on the model training process. Figure 5(a) illustrates the distribution of Reynolds numbers in the test set, which are uniformly distributed over a wide range from $2\times 10^{6}$ to $2\times 10^{7}$. Figure 5(b) shows the variation of the experimental loss with the number of training iterations for the backbone network, revealing that the loss stabilizes after approximately $5\times 10^{5}$ iterations. Finally, Figure 5(c) demonstrates that the model’s performance on the test domain becomes stable after more than $2\times 10^{5}$ training iterations.
Figure 5(d) evaluates the model’s performance with varying sampling steps during testing. Notably, we observe that the prediction accuracy stabilizes after more than three sampling steps. At this point, the time required to predict a single instance is reduced to a matter of seconds, allowing our method to perform near-real-time predictions on the test set.
@@
Training resource consumption. All experiments were conducted on an NVIDIA RTX 6000 GPU. Table 2 compares different model types of DiT (DiT-XL/2, DiT-L/2, DiT-B/2) in terms of GPU memory usage, total number of parameters and training speed (training steps per second). As the model size decreases (from DiT-XL/2 to DiT-B/2), the GPU memory usage significantly drops, while the training speed increases.
@@ IV.3 Case Study
Figure 7: Comparison of surface pressure and velocities of Drela AG09 airfoil at $Re=4.04\times 10^{6}$.
@@
Figure 9: Comparison of surface pressure and velocities of AH 63-K-127/24 airfoil at $Re=9.11\times 10^{6}$.
@@
Figure 11: Comparison of surface pressure and velocities of EPPLER 59 airfoil at $Re=1.21\times 10^{7}$.
Finally, the airfoil is EPPLER 59 at $Re=1.21\times 10^{7}$. The flow-field and surface-quantity comparisons are shown in Figures 10 and 11. Here the model performance degrades, especially in surface pressure and transverse velocity. Two factors likely contribute: (i) fewer high-$Re$ cases at this order in the training set, and (ii) pronounced changes in flow physics when moving from $O(10^{6})$ to $O(10^{7})$, including stronger adverse-pressure-gradient effects and earlier separation onset. These observations suggest that adding more high-$Re$ training samples and/or strengthening physics constraints (e.g., RANS-residual losses) could further improve robustness in this regime.
@@ V Discussion
This study demonstrates the significant potential of the Diffusion Transformer in improving airflow simulations for high Reynolds number airfoil flows. We evaluated AeroDiT against traditional CFD results using three distinct airfoil types and Reynolds numbers ranging from 2 million to 20 million, which were not included in the training dataset.
For the Drela AG09 airfoil at a Reynolds number of $4.04\times 10^{6}$, AeroDiT showed strong agreement with CFD, with a maximum absolute error of approximately 0.02 in the pressure field and no more than 0.06 in the velocity field. In a more challenging test case with the AH 63-K-127/24 airfoil at a Reynolds number of $9.11\times 10^{6}$, AeroDiT performed even better, with nearly all errors in the pressure and velocity fields remaining below 0.02. However, for the EPPLER 59 airfoil at a Reynolds number of $1.21\times 10^{7}$, performance was less satisfactory, likely due to the scarcity of high Reynolds number cases in the training dataset and the significant changes in flow field characteristics at this scale.
@@
These findings highlight the effectiveness of DiT in generalizing across different airfoil shapes and Reynolds numbers, particularly in complex turbulence scenarios common in engineering applications. The study underscores the promise of integrating deep learning with traditional CFD methods, offering a novel approach to address long-standing challenges in accurately simulating airfoil flows.
Future directions could involve expanding the training dataset to include a broader range of high Reynolds number scenarios to enhance the model’ robustness and accuracy in more complex flow fields. Further research will also focus on developing domain generalization techniques Ma et al. (2023); Wang et al. (2022) based on varying Reynolds numbers. By incorporating precise Reynolds numbers as conditional inputs, we aim to further improve the predictive performance of the framework.
@@ VI Conclusion
This study introduces AeroDiT, a novel framework that applies Diffusion Transformers to make real-time, accurate flow field predictions. By leveraging DiTs’ ability to learn conditional distributions $p(\bm{x}|\bm{\Psi})$, AeroDiT achieves reliable predictions of pressure and velocity fields, with average L2 errors of 0.1 for pressure ($p$) and 0.03 and 0.05 for velocity components ($u_{x}$, $u_{y}$), respectively. The results demonstrate the model’ ability to generalize across a broad range of Reynolds numbers ($2.04\times 10^{6}$ to $2.00\times 10^{7}$) and diverse airfoil geometries. In test cases involving the Drela AG09 and AH 63-K-127/24 airfoils, AeroDiT shows excellent agreement with CFD benchmarks, with prediction errors for pressure and velocity fields consistently below 0.02, even under challenging conditions.
Notably, AeroDiT exhibits strong performance in near-real-time scenarios, with the sampling steps converging to stable predictions within seconds. The transformer-based architecture significantly reduces computational costs, making high Reynolds number aerodynamic simulations more feasible for real-time applications. However, in cases involving extreme Reynolds numbers ($1.2\times 10^{7}$), the model’ performance shows some limitations, likely due to the scarcity of corresponding training data and the drastic changes in flow field characteristics at these scales.
@@
These findings highlight AeroDiT’ potential to bridge the gap between computational efficiency and predictive accuracy in CFD simulations. Future work should focus on enriching the training dataset with higher Reynolds number scenarios and extending the model to unsteady and three-dimensional flow conditions, as well as exploring more complex physics constraints.

### v4

@@ I Introduction
The deep learning landscape is still rapidly evolving, with newer architectures such as Transformers Vaswani (2017); Khan et al. (2022); Lin et al. (2022) and denoising diffusion probabilistic models (DDPMs) Ho, Jain, and Abbeel (2020a); Yang and Sommer (2023, 2023)-along with scalable diffusion models integrated with Transformers (DiT) Peebles and Xie (2023a), showing superior performance in tasks like image generation. This evolution raises an important question: Can the latest advancements in machine learning enable real-time, accurate predictions of flow fields, effectively addressing the challenges of computational efficiency and accuracy?
@@
Efficient and accurate prediction of high Reynolds number airflow around airfoils is crucial for advancements in aerodynamics research and the aviation industry, as it directly impacts airfoil optimization, optimal flow control, and the design of more efficient aerodynamic systems. The application of deep learning in aerodynamics has evolved progressively from foundational flow field prediction to efficient optimization design. Early research Sekar et al. (2019) established a speed foundation for replacing traditional CFD through a hybrid CNN-MLP architecture (where CNNs extract geometric features and MLPs map flow parameters), enabling real-time flow field prediction, though limited to steady laminar flows. To address data dependency issues, Wu et al.Haizhou et al. (2022) introduced the daGAN stucture, leveraging a two-phase training mechanism (cGAN pre-training followed by dual-generator fine-tuning) to overcome sparse data constraints and enhance generalization under small-sample conditions. Li et al. Kaandorp and Dwight (2020) developed Deep Learning-based Geometric Filtering (DLGF), utilizing autoencoder-based compression of the design space to reduce variable dimensionality from hundreds to tens, thereby circumventing the curse of dimensionality and accelerating optimization workflows. Within this framework, Wang et al. Muchen et al. (2022) integrated a CNN-based Reduced Order Model (ROM) with a fusion strategy of pooling dimensionality reduction and radial basis functions to handle nonlinear flow optimization under large disturbances. Concurrently, Li et al. Li et al. (2022) addressed low-Reynolds-number separation bubbles by integrating the XFOIL solver with a transition-sensitive turbulence model, achieving efficient airfoil design for UAVs through tailored modal decomposition.
@@ II.2 Diffusion Transformer
In recent years, Transformers have achieved remarkable success across various domains, including natural language processing and computer vision Vaswani (2017). Meanwhile, diffusion models, as a class of generative models, have shown significant potential in image generation tasks. The Diffusion Transformer (DiT) combines the strengths of Transformers with a diffusion-based architecture, demonstrating outstanding performance in diverse applications Peebles and Xie (2023a). Despite these advances, the application of DiT models to aerodynamic simulation remains unexplored. To the best of our knowledge, this study represents the first application of a DiT-based model trained on Reynolds-averaged Navier-Stokes (RANS) simulation data, with a particular focus on predicting flow around airfoils. In addition, in line with prior DiT-based work Ovadia et al. (2023), our study addresses the challenge of real-time inference. A detailed description of the methodology is provided in the following section.
@@ III.4 Forward Diffusion Process
In the forward diffusion process, we followed the canonical procedures described in literature Ho, Jain, and Abbeel (2020a); Sohl-Dickstein et al. (2015). As shown in Figure 2, we gradually add noise to the data $\bm{x}_{0}$ (the true flow field) over several time steps given the input conditions $\bm{\Psi}$, transforming it into a noisy sample $\bm{x}_{T}$. The noise schedule $\beta_{t}$ is pre-defined and fixed, increasing with each timestep. At each timestep $t$, the conditional probability for adding noise is given by:
@@ III.5 Reverse Denoising Process and the Loss
The reverse denoising process $q(\bm{x}_{t-1}|\bm{x}_{t},\bm{\Psi})$ follows a Gaussian distribution, with its mean and variance derived asHo, Jain, and Abbeel (2020a):
@@ III.6 Physics Informed Loss
where $\sigma_{c}$, $\sigma_{p}$, and $\sigma_{v}$ denote the mean absolute values of the convective, pressure, and viscous terms, respectively, used for normalization. Here, $\nabla^{2}u_{x}$ and $\nabla^{2}u_{y}$ are computed component-wise (vector Laplacian), and $\mathbf{g}=(g_{x},g_{y})$ denotes the body force per unit mass. In the present simulations, no external force field is applied, and thus $\mathbf{g}=\mathbf{0}$.
@@ IV.1.1 Datasets and Data preparation
In this study, we utilize three datasets and a test set, as described in Thuerey et al. (2020). These datasets are generated using RANS simulations to compute the velocity and pressure distributions of incompressible flows around airfoils. The simulations cover a wide range of Reynolds numbers ($Re=[0.5,5]$ million) and angles of attack ($\pm 22.5^{\circ}$), providing a diverse set of aerodynamic conditions. A total of 1505 unique airfoil shapes were sourced from the UIUC database and combined with randomly sampled freestream conditions to create the training inputs. The training and test sets are strictly foil-profile disjoint, and the 30 airfoils used for testing do not appear in the 1475 geometries used for training and validation. The RANS solutions were computed using the Spalart-Allmaras (SA) turbulence model, a widely used one-equation model for industrial applications. Boundary conditions are directly embedded into the finite-volume formulation without ghost cells, and gradients and Laplacians are discretized using standard Gauss schemes that remain consistent near masked regions.The computations were carried out with OpenFOAM, an open-source CFD software.
All simulation cases share constant freestream density $\rho$ and dynamic viscosity $\mu$, so variations in Reynolds number arise solely from differences in the freestream velocity $U_{\infty}$ and characteristic chord length $c$. Before training, the velocity and pressure fields are scaled using a characteristic velocity $v_{\mathrm{norm}}$, computed from the maximum magnitude of the input velocity components in each sample. Specifically, velocity components are divided by $v_{\mathrm{norm}}$ and pressure by $v_{\mathrm{norm}}^{2}$. All channels are then normalized using fixed or dataset-specific maxima determined from the training data, and the same normalization constants are applied during inference to ensure consistency. This normalization strategy enforces a consistent non-dimensional representation of the flow fields, allowing the model to implicitly learn across different Reynolds numbers without explicit conditioning.
@@
The test set comprises 90 images, which do not appear in the training data and remain unseen during all stages prior to the testing phase. The Reynolds number distribution in the test set spans a wide range, from $2.04\times 10^{6}$ to $2.00\times 10^{7}$], encompassing 30 different airfoil types. All test cases lie within the same high-$\mathrm{Re}$, fully turbulent regime as the training data and therefore obey the same similarity laws, resulting in nearly identical distributions of non-dimensional flow fields. All the datasets used in this study are publicly available TUM-I Repository 1 (2018); TUM-I Repository 2 (2018).
@@ IV.1.2 Configuration
We employ the DiT (Scalable Diffusion Models with Transformers Peebles and Xie (2023a)) model as the backbone for AeroDiT and adopt the same encoder and decoder models as in Ho, Jain, and Abbeel (2020a). We use three configurations: DiT-B/2, DiT-L/2, and DiT-XL/2. The suffix "/2" indicates that the model operates on a latent feature grid that is downsampled by a factor of 2 in spatial resolution, achieved through a patch size of $2\times 2$ during tokenization.
@@
After training, the conditional encoder is retained as the encoder and the spatial decoder as the decoder. These components are then used to train the DiT model for 2,500,000 iterations with a global batch size of 4 for 150 epochs using the Adam optimizer with an initial learning rate of $1\times 10^{-4}$. The training dataset contained 220,000 unique samples, and an exponential moving average (EMA) of the model parameters with a decay factor of 0.999 was maintained during training and used for final evaluation to improve stability and generalization.
@@ IV.2 Results and Discussion
We provide more specific details on the model training process. Figure 5(a) illustrates the distribution of Reynolds numbers and angles of attack in both training and test sets, confirming that all test cases lie within the global parameter range of the training data, with only a small fraction ($\sim$4–10%) near the distribution boundaries. This indicates that the evaluation covers both in-distribution and mildly out-of-distribution scenarios, ensuring a robust assessment of generalization. Figure 5(b) shows the variation of the experimental loss with the number of training iterations for the backbone network, revealing that the loss stabilizes after approximately $5\times 10^{5}$ iterations. Finally, Figure 5(c) demonstrates that the model’s performance on the test domain becomes stable after more than $2\times 10^{5}$ training iterations.
Figure 5(d) evaluates the effect of the number of diffusion sampling steps on the predictive performance of AeroDiT. The plot explicitly shows the error rates of pressure ($p$), streamwise velocity ($u_{x}$), and transverse velocity ($u_{y}$) as functions of the number of diffusion steps, along with the corresponding inference time. We observe that the prediction error decreases sharply within the first three steps and then plateaus, indicating diminishing returns with additional sampling. Beyond this point, accuracy improvements are marginal, while inference time continues to increase linearly. At around three to five steps, AeroDiT achieves near-optimal accuracy while generating predictions within a few seconds, enabling efficient near-real-time inference on the test set.
@@
IV.3 Computation Efficiency Analysis
All experiments were conducted on an NVIDIA RTX 6000 GPU. Tables 3 and 2 together provide a comprehensive overview of the computational efficiency of AeroDiT. Table 3 compares different DiT model variants (DiT-XL/2, DiT-L/2, DiT-B/2) in terms of GPU memory usage, total number of parameters, and training speed. As expected, smaller models (e.g., DiT-B/2) require significantly less GPU memory and achieve higher training throughput, while larger models (e.g., DiT-XL/2) provide greater representational capacity at a higher computational cost.
Complementing these training-time metrics, Table 2 quantifies AeroDiT’s inference performance under different diffusion sampling steps. Even with increased sampling counts, inference remains efficient — requiring on the order of 1–2 s per sample and achieving throughputs of up to $\sim 1$ sample/s. For comparison, a traditional RANS simulation using OpenFOAM typically requires more than 70 s per case under similar conditions TUM-I Repository 1 (2018). This highlights that AeroDiT not only scales efficiently during training but also delivers near-real-time inference performance suitable for rapid aerodynamic evaluation.
Table 2: Latency and throughput of AeroDiT inference under different diffusion sampling steps (single RTX 6000 GPU).
Sampling Steps Latency (s) Throughput (samples/s)
5 $1.05\pm 0.03$ $0.95\pm 0.03$
10 $1.20\pm 0.03$ $0.83\pm 0.02$
20 $1.50\pm 0.04$ $0.66\pm 0.02$
30 $1.80\pm 0.05$ $0.55\pm 0.02$
50 $2.20\pm 0.06$ $0.45\pm 0.01$
@@ IV.4 Case Study
Figure 7: Comparison of surface pressure and velocities of Drela AG09 airfoil at $Re=4.04\times 10^{6}$. The horizontal axis represents the normalized arc-length coordinate (0–1) measured along the airfoil surface from the leading edge to the trailing edge.
@@
Figure 9: Comparison of surface pressure and velocities of AH 63-K-127/24 airfoil at $Re=9.11\times 10^{6}$. The surface coordinate is again shown as a normalized arc length (0–1) along the airfoil profile.
@@
Figure 11: Comparison of surface pressure and velocities of EPPLER 59 airfoil at $Re=1.21\times 10^{7}$.The surface coordinate is expressed as a normalized arc-length parameter ranging from 0 to 1 along the airfoil profile.
Finally, the airfoil is EPPLER 59 at $Re=1.21\times 10^{7}$. The flow-field and surface quantity comparisons are shown in Figures 10 and 11. Here the model performance degrades, especially in surface pressure and transverse velocity. Two factors likely contribute: (i) fewer high-$Re$ cases at this order in the training set, and (ii) pronounced changes in flow physics when moving from $O(10^{6})$ to $O(10^{7})$, including stronger adverse-pressure-gradient effects and earlier separation onset. These observations suggest that adding more high-$Re$ training samples and/or strengthening physics constraints (e.g., RANS-residual losses) could further improve robustness in this regime.
Although $C_{L}$, $C_{D}$, and $C_{m}$ are not explicitly plotted, their values follow directly from the predicted surface pressure and shear distributions, which closely match CFD results. The accurate capture of pressure gradient reversals and near-wall flow features further ensures reliable identification of separation and reattachment, linking field-level accuracy to performance-critical quantities such as lift, drag, and pitching moment. While integral mass and momentum residuals were not explicitly computed, the close agreement in surface pressure and velocity distributions (Figs. 7, 9, and 11) provides indirect evidence of global flow consistency around the airfoil surface. Future work will incorporate explicit global conservation constraints to enhance physical fidelity further.
@@ V Discussion
This study demonstrates the significant potential of the Diffusion Transformer in improving airflow simulations for high Reynolds number airfoil flows. We evaluated AeroDiT against traditional CFD results using three distinct airfoil types and Reynolds numbers.
For the Drela AG09 airfoil at a Reynolds number of $4.04\times 10^{6}$, AeroDiT showed strong agreement with CFD, with a maximum absolute error of approximately 0.02 in the pressure field and no more than 0.06 in the velocity field. In a more challenging test case with the AH 63-K-127/24 airfoil at a Reynolds number of $9.11\times 10^{6}$, AeroDiT performed even better, with nearly all errors in the pressure and velocity fields remaining below 0.02. However, for the EPPLER 59 airfoil at a Reynolds number of $1.21\times 10^{7}$, performance was less satisfactory, likely due to the scarcity of high Reynolds number cases in the training dataset and the significant changes in flow field characteristics at this scale. In addition, we observe that part of the residual error is concentrated in the near-wall regions, which is expected given the current spatial resolution of $128\times 128$. This grid spacing imposes a limit on the model’s ability to resolve steep pressure gradients and velocity shear within the boundary layer. Furthermore, the latent space representation may introduce some smoothing of fine-scale wall features. While these effects do not significantly affect global aerodynamic predictions, they highlight the importance of resolution-aware modeling when targeting wall-resolved accuracy.
@@
U-Nets remain strong, widely used denoisers in diffusion models Ho, Jain, and Abbeel (2020b); Rombach et al. (2022), whereas AeroDiT adopts a DiT backbone, which scales well and captures long-range, multi-scale dependencies Peebles and Xie (2023b); Pereira et al. (2024). We therefore position AeroDiT as complementary to U-Net surrogates rather than a direct replacement.
Future directions could involve expanding the training dataset to include a broader range of high Reynolds number scenarios to enhance the model’s robustness and accuracy in more complex flow fields. Future work will also explore higher spatial resolutions, multi-scale latent encoding, and wall-focused representations to further improve near-wall fidelity, as well as developing domain generalization techniques Ma et al. (2023); Wang et al. (2022) based on varying Reynolds numbers. By incorporating precise Reynolds numbers as conditional inputs, we aim to further improve the predictive performance of the framework.
@@ VI Conclusion
This study introduces AeroDiT, a novel framework that applies Diffusion Transformers to make real-time, accurate flow field predictions. By leveraging DiTs’ ability to learn conditional distributions $p(\bm{x}|\bm{\Psi})$, AeroDiT achieves reliable predictions of pressure and velocity fields, with average $L_{2}$ errors of 0.1 for pressure ($p$) and 0.03 and 0.05 for velocity components ($u_{x}$, $u_{y}$), respectively. The results demonstrate the model’s ability to generalize across a broad range of Reynolds numbers ($2.04\times 10^{6}$ to $2.00\times 10^{7}$) and diverse airfoil geometries. In test cases involving the Drela AG09 and AH 63-K-127/24 airfoils, AeroDiT shows excellent agreement with CFD benchmarks, with prediction errors for pressure and velocity fields consistently below 0.02, even under challenging conditions.
Notably, AeroDiT exhibits strong performance in near-real-time scenarios, with the sampling steps converging to stable predictions within seconds. The transformer-based architecture significantly reduces computational costs, making high Reynolds number aerodynamic simulations more feasible for real-time applications. However, in cases involving extreme Reynolds numbers ($1.2\times 10^{7}$), the model’s performance shows some limitations, likely due to the scarcity of corresponding training data and the drastic changes in flow field characteristics at these scales.
@@
These findings highlight AeroDiT’s potential to bridge the gap between computational efficiency and predictive accuracy in CFD simulations. Future work should focus on enriching the training dataset with higher Reynolds number scenarios and extending the model to unsteady and three-dimensional flow conditions, as well as exploring more complex physics constraints.